# Efficient Post-Selection for General Quantum LDPC Codes

### v2

@@ 1 Introduction
The logical gap (defined as the log-likelihood ratio difference between candidate corrections in distinct logical classes) has emerged as a Bayesian-motivated metric quantifying decoding confidence [15, 16, 17]. Physically, according to the statistical-mechanical mapping of error correction, the logical gap of a surface code in the code-capacity setting is proportional to the zero temperature free energy cost of a domain wall in the mapped random-bond Ising model [18, 19]. Post-selection strategies based on the logical gap proved remarkably effective, enabling about 15-fold reductions in error rates with a relative overhead factor of <2<2 [15], error threshold improvements (up to 50% with full post-selection under code capacity noise) [17], and order-of-magnitude reductions in magic state preparation costs [20, 21]. The logical gap has become the standard approach for confidence-based post-selection in QEC.
@@
The remainder of this paper is organized as follows. Section 2 reviews the concept of decoder soft outputs and their role in quantifying decoding confidence, including a detailed analysis of the logical gap method and its fundamental limitations. Section 3 introduces our cluster-based confidence metrics and a post-selection strategy based on this, as an alternative approach that overcomes these limitations. Section 4 extends the global strategy to a real-time post-selection strategy based on the sliding-window framework. Section 5 provides detailed performance analysis through numerical simulations for both global and real-time strategies across multiple code families. Finally, Section 6 discusses the implications of our results and outlines directions for future work.
2 Soft outputs from decoding
@@
2.1 Logical gap for quantifying decoding confidence
@@
3 Post-selection strategy based on error cluster statistics
@@
3.1 Clustering-based decoders
@@
3.2 Heuristic confidence metrics from cluster statistics
@@
3.3 Global post-selection strategy
Based on the cluster norm fraction metrics in Definitions 1 and 2, we propose the following post-selection strategy (the term “global decoding” is used to distinguish from the real-time strategy described in Sec. 4):
@@ 3.3 Global post-selection strategy
4 Real-time post-selection strategy
@@
In this section, we extend our post-selection strategy to a real-time setting by integrating it with the sliding-window decoding framework for general QLDPC codes [41, 42, 43, 44, 31, 45]. Rather than collecting clusters statistics from global decoding outcomes, we rely only on data from a constant number of recent windows as shown in Fig. 1(c), enabling us to compute a cluster-confidence metric in real time and to triger early aborts when necessary.
4.1 Sliding-window decoding
The sliding-window decoding framework decomposes the overall decoding problem into smaller, overlapping sub-problems defined on “windows”, which can be solved incrementally. By partitioning along the time direction, a partial correction can be derived from the syndrome data collected so far, enabling real-time decoding. The algorithm is characterized by two positive integers $(W,F)$: the window size $W$ (which governs decoding speed and needs to be small enough to avoid the backlog problem) and the commit size $F$.
The framework works as follows: The set of detectors is partitioned into a sequence of temporal windows, each spanning $W$ rounds, such that two consecutive windows overlap by WFW-F rounds. Within each window, a decoder identifies a correction; however, only a fraction of the correction associated with the first $F$ rounds of each window are committed to the global solution. The syndrome is updated accordingly, and the decoder proceeds to the next window. This process repeats until all detector rounds have been processed.
where $H[i,j]$ denotes the $(i,j)$ element of $H$.
4.2 Post-selection strategy
@@
5 Performance analysis
We present a comprehensive numerical evaluation of the post-selection strategies introduced in Strategies 1 and 2. Our simulations employ the BP+LSD-0 decoder [31] for collecting cluster statistics, which is chosen for its broad applicability to arbitrary QLDPC codes, its error-correcting performance comparable to the BP+OSD decoder [50, 48], and its availability through the open-source Python package ldpc [64]. To ensure cluster statistics are available for all samples, we modified the decoder to execute the LSD subroutine even when BP converges successfully. All simulations use 30 BP iterations with the min-sum method; further optimization of these parameters could potentially enhance the numerical results. Details on our simulation methods are presented in Appendix A.
5.1 Global strategy analysis
@@
We evaluate our approach on a $(3,4)$-regular HGP code with parameters $[[225,9,6]]$, constructed by taking the product of a 12-variable classical LDPC code with itself, whose check matrix is given in Appendix A. This particular code offers practical advantages with its depth-8 syndrome extraction circuit [45], which we employ in our simulations. Figure 4 shows the results at $p=0.001$, where our cluster-based approach achieves approximately two orders of magnitude improvement in $p_{\mathrm{log}}$ compared to baseline methods, which is consistent with the performance gains observed for BB codes.
Additional simulation results are provided in Appendix B, including comprehensive data for all parameter combinations (Figs. 79), a detailed examination of the influence of the norm order $\alpha$ (Fig. 10), and an evaluation of the alternative “conv-max-conf” setting (Fig. 11), which treats BP-converged samples as having maximum decoding confidence. The conv-max-conf setting avoids the need to modify BP+LSD to enforce LSD execution regardless of BP convergence, but is found to degrade post-selection performance in most cases.
5.2 Real-time strategy analysis
We now analyze the real-time post-selection strategy presented in Sec. 4.2. To fairly assess the impact of “aborting mid-way”, we evaluate the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ instead of $p_{\mathrm{abort}}$, by summing the total number of syndrome extraction rounds elapsed across all samples (including both aborted and accepted runs) and dividing by the number of accepted samples. In other words, $\overline{T}_{\mathrm{accepted}}$ represents the expected time cost required to obtain a single accepted run when immediately retrying after each aborted run. For comparison, global strategies yield $\overline{T}_{\mathrm{accepted}}=T/(1-p_{\mathrm{abort}})$.
@@ 6 Remarks
Acknowledgements
We thank Timo Hillmann, Nicholas Fazio, Dominic Williamson, Mingyu Kang, and Hyukgun Kwon for helpful discussions and comments. This work is supported by the Australian Research Council via the Centre of Excellence in Engineered Quantum Systems (EQUS) Project No. CE170100009, and by the Intelligence Advanced Research Projects Activity (IARPA) through the Entangled Logical Qubits program Cooperative Agreement Number W911NF-23-2-0223.
Appendix A Technical details on the simulation methods
This appendix provides technical details on our numerical simulation methods, supplementing the overview given in Sec. 5.
A.1 Circuit generation and sampling
@@ A.3 Decoder implementation
The original decoder implementation skips LSD postprocessing when BP converges, which makes it impossible to obtain cluster-based metrics in converged cases. To address this limitation, we modified the decoder to execute LSD regardless of BP convergence, which is implemented in a fork [^1] https://github.com/seokhyung-lee/ldpc of the ldpc library. Additionally, we test our strategies under the alternative assumption that BP convergence indicates maximum confidence; these results are presented in Fig. 11.
@@
A.4 Overall pipeline implementation
@@
Appendix B Additional figures
Figure 7: Global post-selection analysis with surface codes for all parameter combinations. All shaded regions represent 95% confidence intervals. (a) Logical error rates $p_{\mathrm{log}}$ are plotted against the abort rates $p_{\mathrm{abort}}$ for four decoding confidence metrics: the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$, logical gap, correction weight, and detector density. This is an extension of Fig. 2(b). (b) $p_{\mathrm{log}}$ and $p_{\mathrm{abort}}$ are separately plotted against the cutoff values for the metric $Q_{\mathrm{LLR}}^{(2)}$.
Figure 8: Global post-selection analysis with bivariate bicycle codes for all parameter combinations. All shaded regions represent 95% confidence intervals. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rates $p_{\mathrm{abort}}$ for three decoding confidence metrics: the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$, correction weight, and detector density. This is an extension of Fig. 3(b). (b) $p_{\mathrm{log}}$ and $p_{\mathrm{abort}}$ are separately plotted against the cutoff values for the metric $Q_{\mathrm{LLR}}^{(2)}$.
Figure 9: Global post-selection analysis with the $[[225,9,6]]$ hypergraph product code. $p_{\mathrm{log}}$ and $p_{\mathrm{abort}}$ are plotted against the cutoff values for the metric $Q_{\mathrm{LLR}}^{(2)}$ at $p=0.001$.
Figure 10: Dependence of global post-selection performance on the norm order $\alpha$. Abort rates $p_{\mathrm{abort}}$ required to achieve logical error rate suppression of $10^{-1}$ or $10^{-2}$ (relative to cases without post-selection) are presented for different $\alpha$ values in cluster LLR norm fractions $Q_{\mathrm{LLR}}^{(\alpha)}$. (a), (b), and (c) are for surface, $B$, and HGP codes, respectively. All error bars represent 95% confidence intervals and solid lines are drawn just for visual aid. Overall, $\alpha=1$ and $\alpha=2$ show the best results, while $\alpha=0.5$ and $\alpha=\infty$ have degraded performance.
Figure 11: Global post-selection analysis assuming BP convergence as maximum decoding confidence. Our simulations employ the BP+LSD decoder modified to execute LSD even if BP converges, which is necessary to ensure cluster-based metrics are always obtainable. These plots analyze an alternative approach: assuming converged cases to have maximum decoding confidence (which makes this modification of the decoder unnecessary). $p_{\mathrm{log}}$ is plotted against $p_{\mathrm{abort}}$ for this approach as dashed lines (labeled as “conv-max-conf”) for surface and $B$ codes at $p=0.003$. Since samples with converged BP are always accepted, $p_{\mathrm{abort}}$ is upper bounded by 1(convergence rate)1-$\text${(convergence rate)}. Compared to the original results (solid lines), this alternative approach has degraded performance in most cases, where achievable $p_{\mathrm{log}}$ is lower-bounded by the logical error rate conditioned on convergence.

### v3

# thanks: corresponding authorthanks: corresponding author

@@ 1 Introduction
The logical gap (defined as the log-likelihood ratio difference between candidate corrections in distinct logical classes) has emerged as a Bayesian-motivated metric quantifying decoding confidence [15, 16, 17, 18]. Physically, according to the statistical-mechanical mapping of error correction, the logical gap of a surface code in the code-capacity setting is proportional to the zero temperature free energy cost of a domain wall in the mapped random-bond Ising model [19, 20]. Post-selection strategies based on the logical gap proved remarkably effective, enabling about 15-fold reductions in error rates with a relative overhead factor of <2<2 [15], error threshold improvements (up to 50% with full post-selection under code capacity noise) [17], and order-of-magnitude reductions in magic state preparation costs [21, 22]. The logical gap has become the standard approach for confidence-based post-selection in QEC.
@@
2 Results
2.1 Soft outputs from decoding
@@
2.1.1 Logical gap for quantifying decoding confidence
@@
2.2 Post-selection strategy based on error cluster statistics
@@
2.2.1 Clustering-based decoders
@@
2.2.2 Heuristic confidence metrics from cluster statistics
@@
2.2.3 Global post-selection strategy
Based on the cluster norm fraction metrics in Definitions 1 and 2, we propose the following post-selection strategy (the term “global decoding” is used to distinguish from the real-time strategy):
@@ 2.2.3 Global post-selection strategy
2.3 Real-time post-selection strategy
@@
We extend our post-selection strategy to a real-time setting by integrating it with the sliding-window decoding framework for general QLDPC codes [42, 43, 44, 45, 32, 46]. Rather than collecting clusters statistics from global decoding outcomes, we rely only on data from a constant number of recent windows as shown in Fig. 1(c), enabling us to compute a cluster-confidence metric in real time and to triger early aborts when necessary.
The sliding window decoding framework decomposes the overall decoding problem into smaller, overlapping sub-problems defined on “windows”, which can be solved incrementally. By partitioning along the time direction, a partial correction can be derived from the syndrome data collected so far, enabling real-time decoding. The algorithm is characterized by two positive integers $(W,F)$: the window size $W$ (which governs decoding speed and needs to be small enough to avoid the backlog problem) and the commit size $F$ (which corresponds to the width of overlapping regions between adjacent windows). See the Methods section for more details on the algorithm.
@@
    Here, $\mathcal{E}_{w}^{\text{commit}}$ is defined as the set of error mechanisms committed from window $w$; see Eq. (2) in the Methods section for the formal definition.
@@
2.4 Performance analysis
We present a comprehensive numerical evaluation of the post-selection strategies introduced in Strategies 1 and 2. Our simulations employ the BP+LSD-0 decoder [32] for collecting cluster statistics, which is chosen for its broad applicability to arbitrary QLDPC codes, its error-correcting performance comparable to the BP+OSD decoder [51, 49], and its availability through the open-source Python package ldpc [65]. To ensure cluster statistics are available for all samples, we modified the decoder to execute the LSD subroutine even when BP converges successfully. All simulations use 30 BP iterations with the min-sum method; further optimization of these parameters could potentially enhance the numerical results. Details on our simulation methods are presented in the Methods section.
2.4.1 Global strategy analysis
@@
We evaluate our approach on a $(3,4)$-regular HGP code with parameters $[[225,9,6]]$, constructed by taking the product of a 12-variable classical LDPC code with itself, whose check matrix is given in the Methods section. This particular code offers practical advantages with its depth-8 syndrome extraction circuit [46], which we employ in our simulations. Figure 4 shows the results at $p=0.001$, where our cluster-based approach achieves approximately two orders of magnitude improvement in $p_{\mathrm{log}}$ compared to baseline methods, which is consistent with the performance gains observed for BB codes.
Additional simulation results are provided in the Supplementary Material, including comprehensive data for all parameter combinations (Supplementary Figures 1–3), a detailed examination of the influence of the norm order $\alpha$ (Supplementary Figure 4), and an evaluation of the alternative “conv-max-conf” setting, which treats BP-converged samples as having maximum decoding confidence (Supplementary Figure 5). The conv-max-conf setting avoids the need to modify BP+LSD to enforce LSD execution regardless of BP convergence, but is found to degrade post-selection performance in most cases.
2.4.2 Real-time strategy analysis
We now analyze the real-time post-selection strategy. To fairly assess the impact of “aborting mid-way”, we evaluate the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ instead of $p_{\mathrm{abort}}$, by summing the total number of syndrome extraction rounds elapsed across all samples (including both aborted and accepted runs) and dividing by the number of accepted samples. In other words, $\overline{T}_{\mathrm{accepted}}$ represents the expected time cost required to obtain a single accepted run when immediately retrying after each aborted run. For comparison, global strategies yield $\overline{T}_{\mathrm{accepted}}=T/(1-p_{\mathrm{abort}})$.
@@ 3 Discussion
4 Methods
4.1 Sliding-window decoding
We here describe the sliding-window decoding framework that we use for our real-time post-selection. The framework with the window size $W$ and the commit size $F$ works as follows: The set of detectors is partitioned into a sequence of temporal windows, each spanning $W$ rounds, such that two consecutive windows overlap by WFW-F rounds. Within each window, a decoder identifies a correction; however, only a fraction of the correction associated with the first $F$ rounds of each window are committed to the global solution. The syndrome is updated accordingly, and the decoder proceeds to the next window. This process repeats until all detector rounds have been processed.
where $H[i,j]$ denotes the $(i,j)$ element of $H$.
4.2 Technical details on the simulation methods
We here provide technical details on our numerical simulation methods used to obtain the data in the Results section.
4.2.1 Circuit generation and sampling
@@ 4.2.3 Decoder implementation
The original decoder implementation skips LSD postprocessing when BP converges, which makes it impossible to obtain cluster-based metrics in converged cases. To address this limitation, we modified the decoder to execute LSD regardless of BP convergence, and this modification has been integrated into the official ldpc library. Additionally, we test our strategies under the alternative assumption that BP convergence indicates maximum confidence; these results are presented in the supplementary information.
@@
4.2.4 Overall pipeline implementation
@@
Data availability
The aggregated data that support the findings of this study are openly available in Harvard Dataverse at DOI:10.7910/DVN/MNYLTA. The raw data are not uploaded due to their size but are available upon reasonable request to the authors.
Code availability
The source codes used for the simulations in this work, including the ldpc-post-selection package, are available in GitHub: https://github.com/seokhyung-lee/ldpc-post-selection.
Acknowledgements
We thank Timo Hillmann, Nicholas Fazio, Dominic Williamson, Mingyu Kang, and Hyukgun Kwon for helpful discussions and comments. We also thank Joschka Roffe for integrating our BP+LSD customization into the official ldpc library.
Funding
This work is supported by the Australian Research Council via the Centre of Excellence in Engineered Quantum Systems (EQUS) Project No. CE170100009, and by the Intelligence Advanced Research Projects Activity (IARPA) through the Entangled Logical Qubits program Cooperative Agreement Number W911NF-23-2-0223.
Author contributions
S.H.L. conceived the idea and all authors contributed to developing it. S.H.L. conducted simulations and all authors interpreted the results together. S.H.L. mainly wrote the manuscript and L.E. made the first draft of Figure 1. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.